If sam is twice as old as harry and the sum of their age is the same as 5 x Harrys age decreased by 28 . How old is each ?
Found 2 solutions by stanbon, Jstrasner:
Answer by stanbon(75887)

If sam is twice as old as harry and the sum of their age is the same as 5 x Harrys age decreased by 28 . How old is each ?
=====
Equations:
s = 2h
s+h = 5h-28
---
Substitute for "s" and solve for "h":
2h + h = 5h-28
3h = 5h-28
2h = 28
h = 14 (Harry's age now)
----
Solve for "s":
s = 2h
s = 28 (Sam's age now)
============================
Cheers,
Stan H.
===========

Answer by Jstrasner(112)

Hey,
So for this one, we need to put this question into number form:
S = 2H
S + H = 5(H-28)
Then we substitute 2H for S in the second equation:
2H + H = 5(H-28)
Then we distribute 5 over the right side:
3H = 5H - 140
then we add 140 to both sides and subtract 3H from both sides:
3H - 3H + 140 = 5H - 3H - 140 + 140 =>
140 = 2H
Then we divide both sides by 2:
H = 70
I hope this helps!